OpenStudy (anonymous):

Is there somebody who cen tell me how to multiply this step by step? I am not sure about that! ax(ax+x^2)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ax(ax+x^2)\\~\\ =ax\times ax+ax\times x^2\\~\\ =a^2x^2+ax^3}\end{align}\)

OpenStudy (anonymous):

Okay, I was right, thank you! :)